Width of a tree is defined as the largest distance(number of edges in the shortest path) between any two vertices. What is the smallest possible width of the tree with 100 vertices?

1


2


99


101
1 point
How many complete graphs with at least 1 edge and at most 50 edges are there?

8


9


10


11
1
Expert's answer
2020-11-05T17:27:32-0500

(1) Consider a graph called star. This graph is a special kind of tree with one internal node and k leaves and has the smallest possible width. A star "S_{99}" has 99 edges and 100 vertices and its width is 2.

Answer: 2.


(2) The complete graph "K_n" on n vertices has "n(n-1)\over2". In order to answer that question, we need to solve equation for positive integer n:


"1\le {n(n-1)\over2} \le 50""2\le n^2 - n \le 100"

"n \in \{2, 3,4,5,6,7,8,9,10\}"

Therefore, there are 9 complete graphs with at least 1 edge and at most 50 edges ("K_2, K_3, \dots, K_{10}").

Answer: 9 graphs.